Let W {(x, y, z) R³ : x + y + z = 0}. Check
if W is a subspace of R³. Find a non-zero
subspace U of R3 so that W(intersection)U = (0)
1
Expert's answer
2020-12-17T18:43:31-0500

Let u = (x1,y1,z1) and v = (x2,y2,z2) be two vectors from W. Then:

(1) u+v = (x1+x2, y1+y2, z1+z2) belongs to W, as

(x1+x2) + (y1+y2) + (z1+z2) = (x1 + y1 + z1) + (x2 + y2 + z2) = 0 + 0 = 0

(2) for any constant a the vector au = (ax1,ay1,az1) belongs to W, as

(ax1+ ay1+az1) = a(x1 + y1 + z1) = a*0 = 0

With (1) and (2) we have proved that W is a vector subspace of R³.


Let V = {(x, y, z) R³ : x = y = 0}. Lets prove that V is a subspace of R³.

Let u = (x1,y1,z1) and v = (x2,y2,z2) be two vectors from V. Then:

(1) u+v = (x1+x2, y1+y2, z1+z2) belongs to V, as

x1+ x2 = 0 + 0 = 0 and y1+ y2 = 0 + 0 = 0

(2) for any constant a the vector au = (ax1,ay1,az1) belongs to V, as

ax1= a*0 = 0 and ay1= a*0 = 0.

With (1) and (2) we have proved that V is a vector subspace of R³.


"W \cap V = {0}" , because of all vectors from this set satisfies the system of equations:

x + y + z = 0

x = 0

y = 0

It is clearly, that only (0, 0 , 0) satisfies this system.